OpenStudy (anonymous):

PLEASE HELP WITH THIS CALCULUS PROBLEM! Suppose that you deposit $1.00 into an account that pays 100% annually. a) How much will you have in the bank at the end of 1 year? SOLVED: $2.00 b) Suppose that the interest is compounded twice each year. How much will you have in the bank at the end of 1 year? SOLVED: $2.25 c) Calculate the amount in the bank at the end of 1 year with more compounding periods: 3, 4, 100. d) What number does the amount in part c) seem to be approaching, as the number of compounding periods increases. SHOW ALL WORK PLEASE!

OpenStudy (anonymous):

Can I use that formula in Calculus, because I dont see it anywhere in my textbook

OpenStudy (anonymous):

\[ A=A_0 \left (1+ \frac r n \right)^n \] Where n is the number of times it is compounded each year and e is the annual interest rate.

OpenStudy (anonymous):

For a great explanation, please refer to this thread. http://openstudy.com/study#/updates/4fa7037ee4b029e9dc366b88

OpenStudy (anonymous):

for parts a, b, and c use A=A0(1+rn)n. for part d that the limit as n->\[\infty\]

OpenStudy (anonymous):

what should I subsititue for part a and b then?

OpenStudy (anonymous):

r not e in my post above

OpenStudy (anonymous):

actually thats the limit as *r* goes to infinity my b

OpenStudy (anonymous):

so for part a, r and n are both 1 right? but but what about part b?

OpenStudy (anonymous):

You keep r the same, but you divide it by n the number of time you are compounding

OpenStudy (anonymous):

so n would be 2?

OpenStudy (anonymous):

what about P?

OpenStudy (anonymous):

The limit when n is tending to Infinity of \[ \left ( 1 + \frac 1 n\right)^n \] is equal to e when is approaching Infinity. This problem is to make you believe that statement above, by asking you to compute it for n=1,2, 5, 100, etc

OpenStudy (anonymous):

when n is approaching infinity

OpenStudy (anonymous):

I dont understand how to do part c)

OpenStudy (anonymous):

for parts a and b i did (1+1/1)^1 and (1+1/2)^2

OpenStudy (anonymous):

Compute \[ \left ( 1 + \frac 1 n \right)^n \] for n=3 then for n=4 and finally for n =100

OpenStudy (anonymous):

Ok thank you

OpenStudy (anonymous):

Ooooh, Jobo. I'm so glad you struggled so hard to get the same answer out of this guy that Alex and I gave you a half hour ago.

OpenStudy (anonymous):

For n=50 \[\left(1+\frac{1}{50.}\right)^{ 50}=2.69159 \]

OpenStudy (anonymous):

k thanks :)

OpenStudy (anonymous):

\[\left(1+\frac{1}{10000.}\right) ^{10000}=2.71815 \] and \[ e\approx2.718282 \]

OpenStudy (anonymous):

it approaches the constant e right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

THANKS :)

OpenStudy (anonymous):

yw